19 Roofing: A flat roof 28 ft long and 16 ft wide has a 3-in depth of water sitting on it. What is the weight of the water on the roof?
Answer by bucky(2189)

The volume of water on the roof can be found by multiplying the length of the roof (28 ft) by
the width of the roof (16 ft) by the depth of the water in feet. The 3 inch depth translates
to 1/4 ft. This means that the volume (V) of the water is:
.
V = 28 * 16 * (1/4)
.
And since 1/4 times 16 is 4, the volume equation becomes:
.
V = 28 * 4 = 112 cubic feet
.
One cubic ft of water weighs 62.4 lbs (Pocket Reference Third Edition, Thomas J. Glover, Sequoia
Publishing, Inc., Littleton, CO, Nov 2005, pg 638).
.
Therefore, the total weight of the water on the roof is:
.
112 cu ft * 62.4 lbs/cu ft = 6988.8 lbs
.
That's nearly 3.5 tons of water pushing down on that roof!!!
.
Hope this helps you to understand the problem and how to calculate the answer.
.